# Few-sample Variational Inference of Bayesian Neural Networks with Arbitrary Nonlinearities

### v1

@@ 3.2 Our approach
$\chi_{0,i}$ $=\mu_{i},\quad$ $\beta_{0,i}$ $=\frac{\kappa}{\kappa+1}$
$\chi_{1,i}$ $=\mu_{i}-\sigma_{i}\sqrt{\kappa+1},\quad$ $\beta_{1,i}$ $=\frac{1}{2(\kappa+1)}$
$\chi_{2,i}$ $=\mu_{i}+\sigma_{i}\sqrt{\kappa+1},\quad$ $\beta_{2,i}$ $=\frac{1}{2(\kappa+1)}$ (5)
@@ 4.1 Training details
All networks were trained using the AdamW optimizer [12] with a learning rate $\alpha=0.01$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$, and a weight decay $\lambda=0.01$. We used Bayes by backprop [6] with the Normal distribution likelihood for the likelihood term in the ELBO loss. We scaled the KL-loss term of the ELBO at each epoch $l=1,2,\ldots,M$ by a multiplicative factor $\phi(l)=\frac{2^{M-l}}{2^{M}-1}$ as in [9]. All results shown were averaged over 10 networks trained identically from randomly seeded initializations. We trained all networks using purely synthetic data generated on-the-fly, so networks never saw the same data more than once during training.

### v4

@@ 3.2 Our approach
$\chi_{0,i}$ $=\mu_{i},\quad$ $\gamma_{0,i}$ $=\frac{\kappa}{\kappa+1}$
$\chi_{1,i}$ $=\mu_{i}-\sigma_{i}\sqrt{\kappa+1},\quad$ $\gamma_{1,i}$ $=\frac{1}{2(\kappa+1)}$
$\chi_{2,i}$ $=\mu_{i}+\sigma_{i}\sqrt{\kappa+1},\quad$ $\gamma_{2,i}$ $=\frac{1}{2(\kappa+1)}$ (5)
@@ 4.1 Training details
All networks were trained using the AdamW optimizer [12] with a learning rate $\alpha=0.001$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$, and a weight decay $\lambda=0.01$. We used Bayes by backprop [6] with the Normal distribution likelihood for the likelihood term in the ELBO loss. We scaled the KL-loss term of the ELBO at each epoch $l=1,2,\ldots,M$ by a multiplicative factor $\phi(l)=\frac{2^{M-l}}{2^{M}-1}$ as in [9]. All results shown were averaged over 10 networks trained identically from randomly seeded initializations. We trained all networks using purely synthetic data generated on-the-fly, so networks never saw the same data more than once during training.
@@ 4.3.4 Results
To verify that our physics-informed activations defined in Section 4.3.2 meaningfully improve network performance, we trained several object localization networks as in Section 4.3.3 but with fixed-length datasets defined prior to training. As shown in Figure 6, our physics-informed activations improve network generalization as defined by the negative log-likelihood over an unseen validation set in the data-starved ($\leq 2^{9}$ training pairs) regime.
@@ Appendix A Supplemental figures
Figure 6: Custom activations improve network generalization. Minimum negative log-likelihood achieved during training over an unseen evaluation set when using no output node activation (black dots) versus the physics-informed activations described in Section 4.3.2 (green x’s). Points represent the mean over 10 models trained independently on the same data starting from different random initializations, while error bars represent the standard deviation over the 10 models.